OpenStudy (clara1223):

find the limit as x approaches 9pi/4 of (cos(x)-1)/6x

OpenStudy (clara1223):

\[\lim_{x \rightarrow \frac{ 9\pi }{ 4 }}\frac{ \cos(x)-1 }{ 6x }\]

OpenStudy (anonymous):

substitute

OpenStudy (thomas5267):

\[ \lim_{x\to0}\frac{\cos(x)-1}{x}=0 \]

OpenStudy (zzr0ck3r):

it is defined on a continuous function :)

OpenStudy (freckles):

what is that one commercial that says plug it in plug it in

OpenStudy (clara1223):

@freckles @satellite73 Alright plugging it in I get \[\frac{ \cos(\frac{ 9\pi }{ 4 })-1 }{ 6\frac{ 9\pi }{ 4 } }\] \[\frac{ \frac{ \sqrt{2} }{ 2 }-\frac{ 2 }{ 2 } }{ \frac{ 54\pi }{ 4 } }\] \[\frac{ \sqrt{2}-2 }{ 2 }\times \frac{ 4 }{ 54\pi }\] \[\frac{ 4(\sqrt{2}-2) }{ 108\pi }\] \[\frac{ \sqrt{2}-2 }{ 27\pi }\] Is there any way to simplify it further?

OpenStudy (clara1223):

is this the final answer?

myininaya (myininaya):

I don't see anything wrong with your answer

myininaya (myininaya):

looks totally awesome!

OpenStudy (clara1223):

Great! thanks so much!